[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jul 30 02:45:27 UTC 2017


HSU Shift Log for Sat Jul 29 2017 22:45:21 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: We now get the ECMWF 0.25deg resolution every 3 hours  
to day 6-- check it out if you haven't seen it.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170730/c7368d52/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list